TUGBOAT

Volume 6, Number 1 ^{/} March 1985

2 Addresses of Officers, Authors and Others 3 Official Announcements

General Delivery 4 Message from the President ^{/} Pierre MacKay 4 ANSI X3V1.8 Liaison Report to the TEX Users Group ^{/} Lawrence A. Beck 5 Acknowledgement of Contributions 6 First principles of typographic design for document production: Corrigenda ^{/}

Richard Southall

Low TEX 7 Editor’s introduction ^{/} Maureen Eppstein

Fonts 8 Generic font file format 12 A bit of doggerel ^{/} G. K. M. Tobin

Output Devices 13 Chart: Output devices and computers 15 Device drivers included on the UNIX TEX distribution tape ^{/} Richard Furuta 15 Index to sample output from various devices 15 A TEX82 spooler for VT and dot matrix printers ^{/} M. L. Luvisetto

and E. Ugolini

Site Reports 16 How to get the latest news via TEXhax · Prime 17 Prime site report ^{/} John Crawford · UNIX 18 Unix TEX site report ^{/} Richard Furuta

· VAX/VMS 20 VAX/VMS site report ^{/} Barry Smith 21 TEX for tourists ^{/} James A. Cumiskey

“small” TEX 26 Table of “small” TEX implementations ^{/} Lance Carnes 27 TEX now on microcomputers ^{/} Addison-Wesley Publishing Company 28 PC TEX ^{/} Lance Carnes

Macros 29 Macros for two-column format ^{/} Craig Platt 31 Trees in TEX ^{/} David Eppstein 36 Recipes and fractions ^{/} Donald E. Knuth

News & Announcements 38 TEX, the program 39 Calendar 39 Conference: TEX for scientific documentation, Como, Italy, May 16–17, 1985 39 TUG Annual Meeting, Stanford University, August 14–16, 1985 40 TEX and reports available from Stanford

Computer Science Department

Advertisements 41 Textset, Inc. — TEX and the IBM PC, TEX-to-PostScript 42 Addison-Wesley — MicroTEX 43 TUG — TEX Lectures on Tape 44 Textset, Inc. — Professional TEXware 445 Talaris 446 Personal TEX, Inc. — PC TEX

cover 3 Intergraph Corporation

Forms 47 Membership application and order form 49 Order form for American Mathematical Society publications 51 TEX82 order form

Supplements The TEXbook: Errata and Changes TEX Users Group Membership List, including Profile of TEX installations

Volume 6, Number 2 ^{/} July 1985

54 Addresses of Officers, Authors and Others 55 Official Announcements

General Delivery 56 TUG 1985 Annual Meeting, Stanford University, August 12–14, 1985

Software 56 Editor’s message ^{/} Helmut Jürgensen 57 A multilingual TÊX ^{/} Michael J. Ferguson

Fonts 59 Mathematical symbols and cyrillic fonts ready for distribution ^{/} Barbara Beeton

Output Devices 64 Chart: Output devices and computers 66 Index to sample output from various devices 66 Graphics commands for TEX — Discussion in TEXhax conference ^{/} Alan Spragens 69 Miscellaneous activity at Texas A&M ^{/} Norman W. Naugle

and Tomas G. Rokicki

Site Reports 73 TEXhax activity at Spring DECUS

· IBM Group 73 CMS TEX site report ^{/} Alan Spragens

· DEC 20 76 Report on experience with TEX80 ^{/} Hans Riesel · HP 9000 ser. 500 8080 A TEX82 implementation on the HP9000 series 500 ^{/} Gregory Marriott

· UNIX 81 Unix TEX site report ^{/} Richard Furuta · VAX/VMS 81 VAX/VMS site report ^{/} Barry Smith

“small” TEX 82 Table of “small” TEX implementations ^{/} Lance Carnes

Macros 83 Some diagonal line hacks ^{/} Amy Hendrickson 86 A (possibly) totally useless macro ^{/} Amy Hendrickson 87 Multi-column layout in TEX80 ^{/} August Mohr

Letters 96 Author vs. proofreader ^{/} Jean-Luc Bonnetain 96 Transatlantic TEX ^{/} John Greene

News & Announcements 97 Calendar

98 International Conference on Text Processing and Document Manipulation,

Nottingham, England, April 14–16, 1986

98 PROTEXT II, Dublin, Ireland, October 21–25, 1985

TUG Business 99 TUG 1984 financial report with comparisons ^{/} Samuel B. Whidden

Advertisements 100 The Metafoundry — Introducing mf medley

101 TEX and reports available from the

Stanford Computer Science Department

102 Addison-Wesley — MicroTEX 103 Textset, Inc. 104 Personal TEX, Inc. — PC TEX is here

cover 3 TEX lectures on tape

Forms 105 Membership application and order form 107 TEX82 order form

Supplements The TEXbook: Errata and Changes TUG Membership List, including Profile of TEX installations

Volume 6, Number 3 ^{/} November 1985

110 Addresses of officers, authors and others 111 Official announcements

General Delivery 112 Message from the outgoing President ^{/} Pierre MacKay 113 From the President ^{/} Bart Childs 114 Statement of principles

Software 115 Packed (PK) font file format ^{/} Tomas Rokicki 121 Hyphenation exception log

Font Forum 122 Editor’s introduction ^{/} Georgia K.M. Tobin

124 Mathematical symbols and cyrillic fonts ready for distribution (revised) ^{/} Barbara Beeton

Output Devices 129 Charts: Output devices and computers 131 Index to sample output from various devices

Site Reports 131 The TEX directories at the “source” — Stanford

132 TEX in the UK and Ireland: TEXline

· Data General 132 Data General distribution news ^{/} Bart Childs

· IBM MVS 133 MVS TEX site report ^{/} Craig Platt · IBM VM/CMS 134 TEX at Washington State University ^{/} Dean Guenther

· Hitac 135 Reports on TEX implementation at HUCC ^{/} Tsunetoshi Hayashi · UNIX 137 Unix TEX site report ^{/} Richard Furuta

· VAX/VMS 139 VAX/VMS site report ^{/} Barry Smith

Typesetting on PCs 139 Table of “small” TEX implementations

140 Assembling a moderately-priced, high-performance clone of the IBM PC

for running TEX ^{/} Mitch Pfeffer and Alan Hoenig

146 Review: The Boston Computer Society’s IBM PC & compatibles technical word processor review ^{/} Barbara Beeton

Warnings & Limitations 150 The PLAIN truth: \buildrel ^{/} Barbara Beeton

Macros 150 Announcement of L^{A}TEX Version 2.09 ^{/} Leslie Lamport 151 A solution to the tower of Hanoi problem using TEX ^{/} Bruce Leban 155 Wiggly lines ^{/} A. G. W. Cameron 156 Stream lists and related list types for L^{A}TEX ^{/} Reinhard Wonneberger 158 L^{A}TEX command summary (Version 2.05) available from TUG

Letters 158 Inquiry: Historical compilations ^{/} Rick Hendricks 158 Inquiry: 35mm slides with SliTEX ^{/} Kathy Hornbach

News & Announcements 159 Calendar 159 Acknowledgement of contributions 160 TEX for Scientific Documentation, Strasbourg, France, June 19–21, 1986

cover 3 TUG Annual Meeting, Tufts University, July 28–31, 1986

TUG Business 161 Proposed TUG 1986 budget ^{/} Samuel B. Whidden

Advertisements 162 Addison-Wesley — MicroTEX 163 Textset, Inc. 164 TUG — TEX lectures on videotape 166 Personal TEX, Inc. — PC TEX

Forms 167 TUG membership application 169 TUG order form 173 TEX82 order form

Supplements The TEX book: Errata and Changes TUG Membership List, including Profile of TEX installations

Volume 7, Number 1 ^{/} March 1986

4 Addresses 6 Calendar

TUG News 7 From the President ^{/} Bart Childs 7 Acknowledgment of contributions 7 Donald E. Knuth Scholarship 8 Special credits for this issue 8 Conference: TEX for Scientific Documentation 8 TEX Courses 9 Software Column: Call for Papers 10 Addenda: The BCS Word Processor Review

Letters 11 L^{A}TEX input to ACM journals ^{/} Leslie Lamport 11 Translators to generate TEX files ^{/} Gordon Howell

Short Reports 12 News from the TEX Project ^{/} David Fuchs 12 CDC Cyber Site Report ^{/} Jim Fox 12 Data General Site Report ^{/} Bart Childs 13 Macintosh Site Report ^{/} Barry Smith 13 UNIX Site Report ^{/} Pierre MacKay 15 VAX/VMS Site Report ^{/} Barry Smith 15 L^{A}TEX News ^{/} Leslie Lamport 16 Multilingual TÊX update ^{/} Michael Ferguson

TEX Notes 17 Automatic page sizing ^{/} Alan Auerbach 17 Where to find TEX file descriptions

Articles 18 The PLAIN truth: \displaylines, \ialign ^{/} B. Beeton 20 Multiple changefiles in WEB ^{/} W. Appelt and K. Horn 22 Error detecting changes to TANGLE ^{/} R. M. Damerell 25 DVItoVDU: A TEX page previewer ^{/} A. Trevorrow 37 MacDraw pictures in TEX documents ^{/} H. Varian and J. Sterken 41 DVILASER/PS extensions to L^{A}TEX ^{/} D. Maus and B. Baker 48 Results of the 1985 TUG Questionnaire ^{/} P. Fina 51 TEX text editors for the IBM PC ^{/} A. Hoenig and M. Pfeffer 55 TEX output previewers ^{/} M. Senn 57 TEX output devices ^{/} B. Beeton 67 Memory Sizes in Various Implementations of TEX ^{/} B. Beeton 71 “Small” TEX implementations ^{/} L. Carnes

TUG Business 73 Institutional Members 88 Index of advertisers

Volume 7, Number 2 ^{/} June 1986

91 Addresses

General Delivery 93 Comments on the format of this issue ^{/} Barbara Beeton 93 Computers & Typesetting coming out party ^{/} Barbara Beeton 93 Introducing Donald Knuth and Computers & Typesetting ^{/} Peter Gordon 95 Remarks to celebrate the publication of Computers & Typesetting ^{/}

Donald Knuth

98 Comments on document design prompted by the new TUGboat format ^{/} Cal Jackson

Software 99 VAX Language Sensitive Editor Templates and Guide for use with L^{A}TEX ^{/} Kathy Hornbach

Font Forum 100 Georgia K.M. Tobin 101 The TEX logo in various fonts ^{/} Donald E. Knuth 102 Meta- : An exhibit at the Cooper Union, NYC ^{/} Alan Hoenig

Output Devices 102 Addenda to the output device charts

Site Reports 103 TEX on the Amiga ^{/} Tomas Rokicki

· CDC Cyber 103 CDC Cyber site report ^{/} Jim Fox · Prime 103 Prime 50 Series site report ^{/} John Crawford

Typesetting on PCs 104 The Sperry IT: An IBM AT compatible ^{/} Mitch Pfeffer and Alan Hoenig

Macros 106 Table Making with INRSTEX ^{/} Michael J. Ferguson

Queries 109 WEB system extensions ^{/} Helmut Becker 109 Drama scripts ^{/} John Kennedy 110 Marking changes in revised documents ^{/} Sylvester Fernandez 110 Side-by-side source/output samples;

First-line special handling ^{/} James Alexander

111 Customized editors for TEX ^{/} Karl Kleine 111 Post-L^{A}TEX Index Formatting ^{/} Jim Ludden

Letters 111 Mike Spivak

News & Announcements 112 Calendar 113 Institutional members

Forms 115 TUG membership application 117 TEX82 order form

Advertisements 119 Addison-Wesley — Computers & Typesetting 120 Personal TEX, Inc. — TEX for the PC 121 Talaris — Talaris 810 122 Computer Composition Corporation — Typesetting services 123 Textset, Inc. — New directions 124 FTL systems Inc. — TEX for the Macintosh

cover 3 Addison-Wesley — TEX news

Supplements Computers & Typesetting: Errata and Changes TUG Membership List – Supplement

Volume 7, Number 3 ^{/} October 1986

127 Addresses

General Delivery 129 Message from the President ^{/} Bart Childs 130 Donald E. Knuth Scholarship 131 1987 Annual Meeting, Call for papers: TEX for the Humanities ^{/} Pierre MacKay 132 Report from ANSI X3V1 ^{/} Lawrence A. Beck 132 Towards a TEX philology group ^{/} Reinhard Wonneberger 133 An idea exchange on TEX training ^{/} Laurie Mann

Software 134 Another approach to multiple changefiles ^{/} Klaus Guntermann and Wolfgang Rülling 134 WEB adapted to C ^{/} Klaus Guntermann and Joachim Schrod 138 TIb: a reference setting package for TEX ^{/} James Alexander 140 Packed file format update ^{/} Tomas Rokicki 145 Hyphenation exception log

Fonts 146 Notes on typeface protection ^{/} Charles Bigelow 151 Building Computer Modern Fonts ^{/} John Sauter

Output Devices 152 TEX and Macintosh – new directions in Preview ^{/} Rick Jansen 159 Crudetype: An adaptable device driver ^{/} R. M. Damerell

Site Reports 164 TEXhax returns ^{/} Malcolm Brown

· Atari 164 Porting TEX to the Atari ST ^{/} Klaus Guntermann · CDC Cyber 164 CDC Cyber site report ^{/} Jim Fox · IBM MVS 165 MVS TEX site report ^{/} Craig Platt

· IBM VM/CMS 167 CMS TEX site report ^{/} Alan Spragens

Typesetting on PCs 167 Real typesetting from your personal computer ^{/} Alan Hoenig and Mitch Pfeffer

Macros 169 Harnessing TEX to compute third root of unity primes ^{/} Klaus Thull 171 Statistical graphics with TEX ^{/} Hans Ehrbar

L^{A}TEX 176 Call for participation ^{/} Jackie Damrau and Ken Yap 176 L^{A}TEX bugs ^{/} Leslie Lamport 177 Chapter mottos and optional semi-parameters ^{/} Reinhard Wonneberger 186 A L^{A}TEX addition for formatting indexes ^{/} Thomas Hofmann 187 IdxTEX and GloTEX – indexes and glossaries ^{/} Richard L. Aurbach

Queries 187 Form letters ^{/} John Lee 187 Change bars ^{/} Jim Fox

Letters 188 Bugs in ware ^{/} Wayne Sewell

News & Announcements 190 Calendar 191 INRIA course: Structures for Documents, Aussois, France, January 19–23, 1987 191 , TEX and the Humanities, Rennes, France 191 Japanese TEX Users Group ^{/} Yoshio Ohno 192 TEX in Europe, Summer 1986 ^{/} Barbara Beeton 193 Contents: Proceedings of “TEX for Scientific Documentation”,

Strasbourg, France, June 19–21, 1986

196 Institutional members

Late-Breaking News 194 The Tiger-TEX Workstation ^{/} Susanne Lachmann

Forms 199 TUG membership application 201 TEX order form

Advertisements 198 Index of advertisers cover 3 TUG — Job opening, Technical assistant

Supplements Computers & Typesetting: Errata and Changes TUG Membership List

Volume 8, Number 1 ^{/} April 1987

3 Addresses

General Delivery 5 President’s message (Bart Childs); Acknowledgement of contributions; Donald E. Knuth Scholarship; It happened: announcement of TEX 2.1 (Donald Knuth); News of TEX users in Mexico (Kurt Bernardo Wolf ); Attention, Australian TUG members (Linda Krick)

8 Book publishing using TEX ^{/} Tony Siegman 12 TEX Output for the Future ^{/} Leslie Lamport

Software 12 WEB adapted to C, another approach ^{/} Silvio Levy 14 Mixing right-to-left texts with left-to-right texts ^{/} Donald Knuth and Pierre MacKay

Font Forum 26 Some empirical observations on design ^{/} Georgia K.M. Tobin 29 Write-white printing engines and tuning fonts with ^{/} Neenie Billawala 33 mode def settings for various TEX output devices ^{/} Barbara Beeton

Typesetting on PCs 34 A new TEX-based book typesetting package for the Macintosh

Output Devices 34 Charts: TEX output devices 41 Public domain TEX DVI driver family ^{/} Nelson H. F. Beebe 42 DVItoVDU 1.7 and PSPRINT 1.1 ^{/} Andrew Trevorrow 45 Benson-Varian 9211 looking for a home

Site Reports 46 Notes from TEXhax ^{/} Malcolm Brown

· Atari 46 TEX&Co. on the ST, Part 2 ^{/} Edgar Fuß · Data General 47 Data General distribution news ^{/} Bart Childs

· HP 1000 47 TEX82 and 84 for the HP1000 A-series ^{/} Tor Lillqvist · Prime 48 Prime 50 series site report ^{/} John Crawford

Macros 49 Multiple, independent marks ^{/} Jim Fox 53 Form letter macros ^{/} John S. Garavelli 54 AutoLetter: A TEX form letter procedure ^{/} Edwin V. Bell, II

L^{A}TEX 58 Contents of L^{A}TEX style collection as of 15th February 1987 ^{/} Ken Yap 60 Form letters ^{/} Graeme McKinstry 62 L^{A}TEX’s Index Processing ^{/} Maurizio Zocchi 63 Typesetting ‘Normaltext’ ^{/} Reinhard Wonneberger

Problems 73 TEX does windows (Alan Hoenig); Problem for a Saturday afternoon (Donald Knuth)

Queries 74 APA style in L^{A}TEX (Jennifer Dyck); Chemical formulae (Judy Hawkins); Ancient non-Roman languages (James R. Celoni, S.J.); International Phonetic Alphabet (Boyd Michailovsky); Line numbering (Lawrence D. Cutter); Dictionary formatting (David Nash); Printing out selected pages (Helen S. Horstman)

Letters 75 Wanted: Help for beginners ^{/} Raymond A. Ryan 76 More hyphenation exceptions ^{/} Anders Thulin 76 New fonts for mathematics? ^{/} William A. McWorter

News & Announcements 77 Workshop on Font Design Systems (Sophia-Antipolis, France, 18–19 May 1987)

77 Special Spring Quarter course (Stanford, Donald Knuth):

TEX: The Program: A case study in software design

78 Call for Papers: Electronic Publishing, Document Manipulation and Typography

(Nice, France, 20–22 April 1988)

79 Calendar 81 Institutional members

Forms 83 TUG membership application 85 TEX order form

Advertisements cover 3 Index of advertisers cover 3 TUG — Job opening, Technical assistant

Supplements Computers & Typesetting: Errata and Changes TUG Membership List

Volume 8, Number 2 ^{/} July 1987

99 Addresses

General Delivery 101 From the President ^{/} Bart Childs 101 How to improve the chances for acceptance of your scientific paper ^{/} Gideon Koren

Software 102 TIb: a reference setting package, update (James Alexander); Portuguese hyphenation table for TEX (Pedro de Rezende); A (hopefully) final extension of multilingual TÊX (Michael Ferguson)

103 Report on JTEX: A Japanese TEX ^{/} Yasuki Saito 117 Multiple changefiles: the adventure continues ^{/} E. Wayne Sewell 118 How to MANGLE your software: the WEB system for Modula-2 ^{/} E. Wayne Sewell 123 The SCANTEX processor ^{/} E. Wayne Sewell

Fonts 128 Blacker Thoughts ^{/} John S. Gourlay 129 Updated Computer Modern fonts for the LN03 ^{/} John Sauter 131 MFtool: A script-driven processing facility ^{/} John Crawford 132 Update: mode def settings for TEX output devices ^{/} Barbara Beeton 135 Fonts for digital halftones ^{/} Donald E. Knuth

Output Devices 161 The ideal TEX driver ^{/} Robert W. McGaffey 163 TEX output devices (with charts) ^{/} Don Hosek 171 Using PostScript with TEX ^{/} Alec Dunn 237 Index to sample output from various devices

Site Reports 174 TEX-L access for Bitnet users ^{/} Glenn L. Vanderburg

174 TEX and training: A case study ^{/} Laurie Mann

· Data General 176 Data General site report ^{/} Bart Childs · IBM VM/CMS 178 IBM VM/CMS site report (Dean Guenther); SAS merged with TEX (Gil Pierson)

· UNIX 179 Unix TEX site report ^{/} Pierre MacKay · VAX/VMS 177 Enhancements to TEX on the VAX ^{/} Adrian F. Clark

Typesetting on PCs 181 A bug in TEXTURES v0.95 prerelease ^{/} Gerhard F. Kohlmayr 181 Work-around for an ImageWriter problem affecting TEXTURES output ^{/} Barry Smith 182 Update: Real typesetting from your PC ^{/} Alan Hoenig and Mitch Pfeffer

Macros 182 Macros with keyword parameters ^{/} Wolfgang Appelt 184 \expandafter in conditionals; a generalization of PLAIN’s \loop ^{/} Alois Kabelschacht 185 TEX in the commercial environment — multi-column output ^{/} Elizabeth Barnhart 190 Diglot typesetting ^{/} Charles LeHardy 193 First line special handling with TEX ^{/} Anne Brüggemann-Klein

L^{A}TEX 198 Contents of L^{A}TEX style collection as of 15th May 1987 ^{/} Ken Yap 201 Automated index generation for L^{A}TEX ^{/} Richard Aurbach

Problems 210 Saturday morning problem — solution (Donald Knuth); TEX does windows — conclusion (Alan Hoenig); Comment on “TEX does windows” (Jim Fox)

Queries 216 Time line macro (James Alexander); Reply: Printing out selected pages (Donald Knuth); Using the Windows environment (Mike Black)

News & Announcements 218 TUG Annual Meeting, University of Washington, August 24–26, 1987 219 Calendar 220 PROTEXT IV (Boston, Mass., 22–24 October 1987); 6th German TEX meeting

(University of Münster, 8–9 October 1987)

Late-Breaking News 221 SGML and TEX ^{/} Lynne A. Price

TUG Business 225 Treasurer’s report ^{/} Samuel B. Whidden 228 Institutional members

Forms 231 TUG membership application; TEX order form

Advertisements 252 Index of advertisers

Supplements Computers & Typesetting: Errata and Changes; TUG Membership List Supplement

Volume 8, Number 3 ^{/} November 1987

255 Addresses

General Delivery 257 From the President ^{/} Bart Childs 258 The Volunteer Tree ^{/} Rilla Thedford 259 Book publishing using TEX ^{/} Charles-Michel Marle 260 TEX training, etc. — A TUG meeting trip report ^{/} Laurie Mann 262 From the Editor ^{/} Barbara Beeton

Software 262 Proposed minimum standards for TEX distributions ^{/} Bart Childs

264 TANGLE modification causes problems in and PK files ^{/}

Thomas J. Reid

266 Hyphenation exception log

Fonts 268 Update: mode def settings for TEX output devices ^{/}

Doug Henderson

270 Halftone output from TEX ^{/} Adrian F. Clark 275 Generating an APL font ^{/} Aarno Hohti and Okko Kanerva

Output Devices 279 TEX output devices (with charts) ^{/} Don Hosek 287 DVI driver considerations for high-volume printing systems ^{/} Thomas J. Reid 291 \special issues ^{/} Glenn L. Vanderburg and Thomas J. Reid

Site Reports 301 TEX information for users in the U.K. ^{/} Graham Toal 301 TEXhax Notes ^{/} Malcolm Brown 304 6th Meeting of the “TEX-Interessenten” in Germany ^{/} Joachim Lammarsch

· Atari ST 304 Atari ST site report ^{/} Klaus Guntermann · Data General 304 Data General site report ^{/} Bart Childs

Typesetting on PCs 305 Running TEX on a 386-based computer: Twice as fast as an AT ^{/} Mitch Pfeffer and Alan Hoenig

Macros 307 What constitutes a well-documented macro? ^{/} Christina Thiele 309 Macros for Jill ^{/} Donald E. Knuth 315 Floating figures at the right, and Some random text for testing ^{/} Thomas J. Reid

L^{A}TEX 320 Contents of L^{A}TEX style collection as of 15th September 1987 ^{/} Ken Yap 323 The L^{A}TEX user’s column ^{/} Jackie Damrau

Queries 306 Request for contributions to a new publication ^{/} Peter Flynn 326 Reply: Printing out selected pages ^{/} Jeffery Boes 326 Title formatting macro wanted ^{/} Stephen C. Lipp

News & Announcements 328 Calendar

329 Exeter University: TEX88, 18–20 July 1988 cover 3 Call for papers: TUG Annual Meeting, Montréal, 22–24 August 1988

TUG Business 331 Institutional members

Forms 333 TUG membership application

Advertisements 352 Index of advertisers

Supplements Computers & Typesetting: Errata and Changes, Supplement TUG Membership List Supplement

Volume 9, Number 1 ^{/} April 1988

3 Addresses

General Delivery 5 From the President ^{/} Bart Childs 5 TUGboat docks at fire station ^{/} Alan Wittbecker 6 Donald E. Knuth Scholarship 6 Notes from the Editor ^{/} Barbara Beeton 8 Document Production: Visual or Logical? ^{/} Leslie Lamport

Software 11 Still another aspect of multiple change files: The PATCH processor ^{/}

Peter Breitenlohner

12 Turkish hyphenations for TEX ^{/} Pierre A. MacKay

Font Forum 15 The ABC’s of special effects ^{/} Georgia K.M. Tobin 19 Blackboard Bold ^{/} Robert Messer 20 Using Greek fonts with TEX ^{/} Silvio Levy

Output Devices 25 TEX output devices (with charts) ^{/} Don Hosek 34 An ASCII previewer for TEX ^{/} Marcus Brown 37 A screen previewer for VM/CMS ^{/} Don Hosek 37 Why TEX should NOT output PostScript — yet ^{/} Shane Dunne 39 Index to sample output from various devices

Site Reports

· Amiga 40 The Commodore Amiga: A magic TEX machine ^{/} Tom Rokicki · Data General 41 Data General site report ^{/} Bart Childs

· Fujitsu 42 Fujitsu ports TEX 1.0 onto M-series · IBM VM/CMS 42 IBM VM/CMS site report ^{/} Dean Guenther

· UNIX 42 Unix TEX site report ^{/} Pierre MacKay

44 TEX to C converter ^{/} Tim Morgan

· VAX/VMS 45 VAX/VMS site report ^{/} David Kellerman

Typesetting on PCs 46 Writers’ tools I: PC spelling and grammar checkers ^{/} Alan Hoenig and Mitch Pfeffer 47 Grapevine reports of inexpensive versions of TEX ^{/} Alan Hoenig and Mitch Pfeffer 48 TurboTEX: A new port in C for Unix and MS-DOS ^{/} Richard J. Kinch

and Jennifer L. Vollbrecht

52 easyTEX ^{/} Ester Crisanti, Alberto Formigoni and Paco La Bruna

Macros 57 A tutorial on \expandafter ^{/} Stephan v. Bechtolsheim 61 Macros for outlining ^{/} James W. Walker 64 A macro writing tool: Generating new definitions ^{/} Amy Hendrickson 65 French in TEX ^{/} Alonzo Gariepy 70 German TEX ^{/} Hubert Partl

L^{A}TEX 73 Contents of L^{A}TEX style collection as of 4th February 1988 ^{/} Ken Yap 76 The L^{A}TEX user’s column ^{/} Jackie Damrau 78 Page layout in L^{A}TEX ^{/} Kent McPherson

Queries 83 Automatic page balancing macros wanted; Replies: Inverted pyramidal titles; Logarithmic time scales ^{/} Frederick H. Bartlett

News & Announcements 86 Calendar 87 TUG 1988 Annual Meeting, Montréal, August 22–24 87 Videotapes of Knuth’s software course based on TEX: The Program

TUG Business 88 Institutional members

Forms 91 TUG membership application 95 TEX order form

Advertisements 112 Index of advertisers

Supplements Computers & Typesetting: Errata and Changes TUG Membership List

Volume 9, Number 2 ^{/} August 1988

115 Addresses

General Delivery 117 From the President ^{/} Bart Childs 117 Los Alamos sets new membership record ^{/} Ray Goucher 117 Donald Knuth awarded Franklin Medal 118 TEX in the publishing environment: A survey of production/commercial users ^{/}

Elizabeth Barnhart

121 Editorial comments ^{/} Barbara Beeton

Software 121 New version(s) of TEX and ^{/} Barbara Beeton 122 64-bit TEX ^{/} Bart Childs 123 Porting TEX to C ^{/} Klaus Lichtenwalder 124 TEX adapted to CWEB ^{/} David Kennedy

Fonts 125 Some useful variations of standard fonts ^{/} Glenn Vanderburg 126 Designing for low-res devices ^{/} Georgia K.M. Tobin 129 TEX fonts and suggested magnifications ^{/} Bart Childs 131 The many faces of TEX: A survey of digital fonts ^{/} Dominik Wujastyk 152 A Punk Meta-Font ^{/} Donald E. Knuth

Output Devices 169 TEX output devices (with charts) ^{/} Don Hosek 178 Why TEX should NOT output PostScript — yet: Addendum ^{/} Shane Dunne 178 ASCII Preview with vuTEX ^{/} Warren Wolfe 151 Index to sample output from various devices

Site Reports

· Cray 181 TEX on the Cray ^{/} Bart Childs · Data General 181 Data General site report ^{/} Bart Childs

· Macintosh 181 Macintosh site report ^{/} Barry Smith

Warnings & Limitations 182 Controlling <ctrl-M>; Ruling the depths ^{/} Barbara Beeton

Macros 183 German TEX, a next step ^{/} Peter Breitenlohner 186 Some problems with the INRSTEX table making macros ^{/} Michael J. Wichura 189 Box plots and scatter plots with TEX macros ^{/} A. J. Van Haagen 193 PICTEX: Macros for drawing PICtures ^{/} Michael J. Wichura

L^{A}TEX 198 The L^{A}TEX user’s column ^{/} Jackie Damrau 200 Contents of L^{A}TEX style collection as of 19th June 1988 ^{/} Ken Yap 203 A note on processing parts with L^{A}TEX ^{/} Stephan v. Bechtolsheim 204 Page layout in L^{A}TEX: Erratum ^{/} Kent McPherson

News & Announcements 205 Calendar 206 GUTenberg meeting report ^{/} Bernard Gaulle

Late-Breaking News 207 Knuth Scholarship winner

TUG Business 208 Institutional members

Forms 211 TUG membership application

Advertisements 226 Index of advertisers

Supplements Computers & Typesetting: Errata and Changes, Supplement M–1 TUG Membership List Supplement

Volume 9, Number 3 ^{/} November 1988

229 Addresses

General Delivery 231 From the President ^{/} Bart Childs 232 Extra! Extra! TUGboat becomes a quarterly 232 Donald Knuth awarded Oxford degree 233 TEX and TUG go international — A trip report ^{/} Barbara Beeton 236 Some typesetting conventions ^{/} Graeme McKinstry

Software 238 Software-ergonomics on the Atari ST ^{/} Klaus Heidrich 239 Hyphenation exception log

Fonts 241 L^{A}TEX fonts and suggested magnifications ^{/} Joachim Schrod 244 Mode def’s please ^{/} Doug Henderson 246 Further faces ^{/} Dominik Wujastyk

Output Devices 251 TEX output devices (with charts) ^{/} Don Hosek 261 DVItoVDU 3.0 and PSPRINT 3.0 ^{/} Peter Abbott and Andrew Trevorrow

Site Reports 263 A UK-based TEX mail archive server ^{/} Peter Abbott

· Atari ST 264 Atari ST site report ^{/} Klaus Guntermann · Data General 265 Data General site report ^{/} Bart Childs · IBM VM/CMS 265 IBM VM/CMS site report ^{/} Dean Guenther

· IBM MVS 266 MVS site report ^{/} Craig Platt · UNIX 267 Unix TEX site report ^{/} Pierre MacKay

Typesetting on PCs 269 Recovering from a hard-disk failure ^{/} Mitch Pfeffer and Alan Hoenig 271 Evaluation of K-talk ^{/} C.G. van der Laan and J.R. Luyten

Training 272 Making paragraphs ^{/} Alan Wittbecker

Macros 276 A tutorial on \futurelet ^{/} Stephan v. Bechtolsheim 279 Compact matrix display ^{/} Thomas Kneser 284 Typesetting chess ^{/} Wolfgang Appelt 287 Equation numbering in plain TEX ^{/} J. E. Pittman 289 Loopy.TeX ^{/} J. E. Pittman 292 A page make-up challenge ^{/} David F. Rogers

L^{A}TEX 294 Contents of L^{A}TEX style collection as of 24 September 1988 ^{/} Michael DeCorte 297 The L^{A}TEX column ^{/} Jackie Damrau 298 A new implementation of the array– and tabular–environments ^{/} Frank Mittelbach

News & Announcements 315 Calendar

316 Dutch TEX Users Group ^{/} C. G. van der Laan cover 3 Call for papers: TUG Tenth Annual Meeting,

Stanford University, 21–23 August 1989

316 Call for papers: RIDT’89 — International Workshop on Raster Imagining and Digital Typography, Lausanne, Switzerland, 12–13 October 1989

Late-Breaking News 317 TEXhax moves north ^{/} Pierre MacKay 317 Production notes ^{/} Barbara Beeton 318 A thank you note 324 TUG’s Tenth

TUG Business 318 Accountant’s review of TUG financial records,

years ended December 31, 1987 and 1986

322 Institutional members

Forms 325 TUG membership application

Advertisements 342 Index of advertisers

Supplements M–1 TUG Membership List Supplement

Volume 10, Number 1 ^{/} April 1989

3 Addresses

General Delivery 5 Donald E. Knuth Scholarship 5 From the President ^{/} Bart Childs 5 Announcing a TUG dingbat competition 6 Editorial comments ^{/} Barbara Beeton 8 TEX would find it difficult … ^{/} Donald Knuth 8 A TEX encounter in Japan ^{/} Barbara Beeton

Software 11 News from the VORTEX project ^{/} Michael Harrison 14 An enhanced TEX-editor interface for VMS ^{/} Adrian Clark 15 The virtual memory management of PubliC TEX ^{/} Klaus Thull 23 Turbo : A new port in C for UNIX and MS-DOS ^{/} Richard Kinch 25 The TEX PostScript software package ^{/} Stephan v. Bechtolsheim

Fonts 28 A handy little font ^{/} Georgia K.M. Tobin 31 Typesetting Concrete Mathematics ^{/} Donald Knuth 36 Outline fonts with ^{/} Doug Henderson 39 Font news ^{/} Dominik Wujastyk

Graphics 39 Computer graphics and TEX — a challenge ^{/} David F. Rogers 44 A portable graphics inclusion ^{/} Bart Childs, Alan Stolleis and Don Berryman

Output Devices 46 TEX output devices (with charts) ^{/} Don Hosek 56 Report from the DVI driver standards committee ^{/} Don Hosek 56 High quality printing of TEX DVI output files in the VAX/VMS environment ^{/}

Marius Broeren and Jan van Knippenberg

Query 58 Output driver for Xerox 4045 on IBM 3090 ^{/} Arturo Puente

Site Reports 59 UKTEX and the Aston archive ^{/} Peter Abbott · Data General 60 DG site report ^{/} Bart Childs

· UNIX 60 UnixTEX site report ^{/} Pierre MacKay · VAX/VMS 62 VAX/VMS site report ^{/} David Kellerman

Typesetting on PCs 62 The land of the free and the near free ^{/} Alan Hoenig

65 AmigaTEX… or How envy was resisted and knowledge found on the road to Ööç ^{/}

Kim Kubik

67 Using TEX and L^{A}TEX with WordPerfect 5.0 ^{/} Michael Modest

Macros 73 A page make-up macro ^{/} Joost Zalmstra and David F. Rogers 82 Extended equation numbering in Plain TEX ^{/} James Nearing 89 APE — A set of TEX macros to format Ada programs ^{/} Sriram Sankar

L^{A}TEX 97 Contents of L^{A}TEX style collection as of 16 January 1989 ^{/} Michael DeCorte

103 “A new implementation of the array– and tabular–environments of L^{A}TEX”

(TUGboat 9#3) — addenda ^{/} Frank Mittelbach

105 Drawing histogram bars inside the L^{A}TEX picture–environment ^{/} Rainer Schöpf 108 Vertical centering for transparencies ^{/} Dezső Nagy 113 Typesetting bridge via L^{A}TEX ^{/} C. G. van der Laan

News & Announcements 116 WOODMAN ’89: Workshop on Object Oriented Document Manipulation,

Rennes, France, 29–31 May 1989

117 Calendar 118 GUTenberg Congrès: TEX et les graphiques, Paris, 16–17 May 1989 118 TEX89: Karlsruhe University, 11–13 September 1989 119 A UK TEX users’ group — Report of a preliminary meeting ^{/} David Osborne

Late-Breaking News 120 Production notes ^{/} Barbara Beeton 121 Coming next issue

TUG Business 122 Institutional members

Forms 125 TUG membership application

Advertisements 142 Index of advertisers

Supplements Computers & Typesetting: Errata and Changes

Volume 10, Number 2 ^{/} July 1989

145 Addresses

General Delivery 147 From the President ^{/} Bart Childs 147 Editorial comments ^{/} Barbara Beeton 150 Réflexions sur le Congrès GUTenberg ^{/} Malcolm Clark 153 International Standards and TEX ^{/} Malcolm Clark 156 Teaching TEX ^{/} Bart Childs

Software 164 HI-TEX cutting & pasting ^{/} Michael Spivak, Michael Ballantyne and Yoke Lee

Fonts 166 Another dingbat idea ^{/} Georgia K.M. Tobin 170 Chess printing via and TEX ^{/} Zalman Rubinstein 173 Guidelines for creating portable code ^{/} Don Hosek

Graphics 177 Integration of TEX and graphics at the Pittsburgh Supercomputing Center ^{/}

Phil Andrews

179 METAPLOT: Machine-independent line graphics for TEX ^{/} Patricia Wilcox

Output Devices 188 TEX output devices ^{/} Don Hosek 188 Report from the DVI driver standards committee ^{/} Tom Reid and Don Hosek

Resources 192 Announcing (belatedly) TEXMAG ^{/} Don Hosek 193 TEXline: A newsletter of the TEX community ^{/} Malcolm Clark 194 UKTEX and the Aston archive ^{/} Peter Abbott 195 The DECUS TEX collection ^{/} M. Edward Nieland 196 The Clarkson archive server — Update as of 1 May 1989 ^{/} Michael DeCorte

Site Reports 201 Data General site report ^{/} Bart Childs 201 Prime 50 series site report ^{/} John Crawford

Typesetting on Personal Computers 202 “Free” TEX software for IBM PCs ^{/} Jon Radel 202 Public domain TEX for the Mac ^{/} Andrew Trevorrow

Tutorials 203 \csname and \string ^{/} Stephan v. Bechtolsheim

Macros 207 DDA methods in TEX ^{/} David Salomon 217 Printing Vietnamese characters by adding diacritical marks ^{/} Brother Eric Vogel 222 A new font selection scheme for TEX macro packages — the basic macros ^{/}

Frank Mittelbach and Rainer Schöpf

L^{A}TEX 239 A bar chart in L^{A}TEX ^{/} Dezső Nagy 241 Producing on-line information files with L^{A}TEX ^{/} Hubert Partl 245 The doc–option ^{/} Frank Mittelbach 274 The autodoc–option ^{/} B. Hamilton Kelly

News & Announcements 285 Calendar 287 Nordic TEX meeting, 12 June 1989, Royal Institute of Technology, Stockholm 287 L^{A}TEX course book in Dutch 287 DANTE — Deutschsprachige Anwendervereinigung TEX ^{/} Joachim Lammarsch 288 Notes on first meeting of UK TEX Users’ Group, 15 March 1989 ^{/} David Osborne 289 Production notes ^{/} Barbara Beeton

Forms 290 TUG membership application

Supplements M–1 TUG Membership List Supplement

Volume 10, Number 3 ^{/} November 1989

309 Addresses

General Delivery 311 From the President ^{/} Bart Childs 312 Message from the new President ^{/} Nelson H. F. Beebe 314 Editorial comments ^{/} Barbara Beeton 316 The first Dutch TEX days ^{/} Victor Eijkhout and Nico Poppelier 319 Answers to the TEXtests ^{/} Bart Childs 324 Crossword ^{/} Technites

Software 325 The new versions of TEX and ^{/} Donald E. Knuth 328 PubliC available ^{/} Klaus Thull 329 A .dvi file processing program ^{/} Stephan von Bechtolsheim 332 Notes on Russian TEX ^{/} Dimitri Vulis 336 Hyphenation exception log ^{/} Barbara Beeton

Fonts 342 Erratum: Typesetting Concrete Mathematics, TUGboat Vol. 10, No. 1 342 TEX and latin alphabet languages ^{/} Yannis Haralambous 346 Fonts and PostScript ^{/} Mike Parker 351 A chess font for TEX ^{/} Jan Eric Larsson 352 The road to Ethiopic TEX ^{/} Abass Andulem 354 Typesetting modern Greek with 128 character codes ^{/} Yannis Haralambous

and Klaus Thull

359 Erratum: Chess printing via and TEX, TUGboat Vol. 10, No. 2

Resources 360 Users’ guide to LaTeX-help ^{/} Max Hailperin 360 TEX-Ed ^{/} Yin Kean 362 The Clarkson archive server — Update as of 22 September 1989 ^{/} Michael DeCorte 365 Coming in January from the American Mathematical Society 366 TEX EURO ^{/} Joachim Lammarsch 367 GUTenberg will distribute MLTEX ^{/} Bernard Gaulle 367 A TEX mailbox in Germany ^{/} Garry Glendown

· Atari ST 368 Towards a complete and comfortable TEX System ^{/} Stephan Lindner and Lutz Birkhahn · VAX/VMS 372 VMS site coordinator’s report ^{/} David Kellerman

Typesetting on Personal Computers 373 TEX-PostScript output on non-PostScript devices ^{/} Alan Hoenig and Mitch Pfeffer 376 Controlling AmigaTEX from CygnusEd ^{/} Garry Glendown

Macros 378 TUGboat authors’ guide ^{/} Ron Whitney and Barbara Beeton 385 Round boxes for plain TEX ^{/} Garry Glendown 387 Printing annotated chess literature in natural notation ^{/} Zalman Rubinstein 390 Bibliographic citations; or variations on the old shell game ^{/} Lincoln Durst 394 Macros for indexing and table-of-contents preparation ^{/} David Salomon

Query 400 Scribe-to-TEX (Terry Bush)

L^{A}TEX 400 Towards L^{A}TEX 2.10 ^{/} Frank Mittelbach and Rainer Schöpf

401 The development of national L^{A}TEX styles ^{/} Johannes Braams,

Victor Eijkhout and Nico Poppelier

407 An environment for multicolumn output ^{/} Frank Mittelbach 416 An extension of the L^{A}TEX theorem environment ^{/} Frank Mittelbach

Abstracts 427 Deutsche Kurzfassungen der TUGboat-Artikel ^{/} Luzia Dietsche

News & Announcements 429 Calendar and Calls for Papers 432 Programs of the recent TUG and EuroTEX Conferences

Late-Breaking News 437 Production notes ^{/} Barbara Beeton

TUG Business 438 Institutional members 440 Treasurer’s report ^{/} David Ness

Forms 444 TUG membership application

Advertisements 461 Index of advertisers

Supplements M–1 TUG Membership List Supplement

Volume 10, Number 4 ^{/} December 1989 1989 Conference Proceedings

Introduction 465 Christina Thiele ^{/} Editor’s introduction

Font Forum 467 Doug Henderson ^{/} Introduction to 481 Neenie Billawala ^{/} Opening Pandora’s box 491 Alan Hoenig ^{/} Fractal images with TEX 499 Don Hosek ^{/} Design of Oriental characters with 503 Bob Batzinger ^{/} Thai languages and 505 John D. Hobby ^{/} A -like system with PostScript output 513 R.E. Youngen, W.B. Woolf and D.C. Latterner ^{/} Migration from Computer

Modern fonts to Times fonts

521 Arvin C. Conrad ^{/} Fine typesetting with TEX using native Autologic fonts

KEYNOTE ADDRESS 529 Donald E. Knuth ^{/} The errors of TEX

TEX Training 533 Michael Doob ^{/} Of the computer scientist, by the computer scientist, for the

computer scientist

541 Hope Hamilton ^{/} Mastering TEX with templates 549 Anita Z. Hoover ^{/} Using WordPerfect 5.0 to create TEX and L^{A}TEX documents 561 Robin L. Kubek ^{/} TEX for the word processing operator 567 Jo Ann Rattey-Hicks ^{/} TEX and its versatility in office production

General Applications 579 Max Díaz ^{/} TEX in México 595 James Haskell, Wally Deschene and Alan Stolleis ^{/} TEX for 30,000 603 Alan E. Wittbecker ^{/} TEX enslaved

Graphics Applications 607 J.T. Renfrow ^{/} Methodologies for preparing and integrating PostScript graphics

627 Rolf Olejniczak-Burkert ^{/} texpic — Design and implementation of a picture graphics language in TEX à la pic

Database Applications 639 D.C. Latterner and W.B. Woolf ^{/} TEX at Mathematical Reviews 655 Jörgen L. Pind ^{/} Lexicography with TEX

General Information 667 Malcolm Clark ^{/} Olde worlde TEX 675 Peter Abbott ^{/} The UKTEX archive at the University of Aston

TEX Tools 681 Frank Mittelbach and Rainer Schöpf ^{/} With L^{A}TEX into the nineties 691 Andrew Marc Greene ^{/} TEXreation — playing games with TEX’s mind 707 Bill Cheswick ^{/} A permuted index for TEX and L^{A}TEX 715 Steve Sydoriak ^{/} L^{A}TEX memos and letters 727 Gary Benson, Debi Erpenbeck and Janet Holmes ^{/} Inserts in a multiple-column

format

743 Mary McClure ^{/} TEX Macros for COBOL syntax diagrams 751 Don L. Riley and Brad L. Halverson ^{/} Creating an efficient and workable PC

interface for TEX

Forms 760 TUG membership application

Participants 761 List of participants at the Stanford meeting (August 20–23, 1989)

Advertisements 765 Index of advertisers